Notifications

Posted: April 3rd, 2020 at 16:00 by Iain Beveridge

Notifications of new posts to noc.enta.net will no longer automatically appear on any of the Entanet twitter feeds. If you wish to receive notifications of new posts, please sign up to email notifications via the link on the right hand side of this page.

Incident: L2TP broadband

Posted: March 31st, 2020 at 11:46 by david labouchardiere

We are currently aware of an issue affecting authentication on L2TP broadband circuits. Engineers are already engaged and are investigating.

Further updates will be provided as and when they become available.

 

  1. david labouchardiere says:

    Engineers confirm service has now been restored. Please ask any customers still having issues to reboot their router before contacting the service desk. We apologise for any inconvenience caused.

Incident: slo24-rtr-1

Posted: March 18th, 2020 at 14:40 by Jordan Coleman
We are currently aware of an issue affecting the slo24-rtr-1 PE. Initial diagnostics indicate a line card failure. Engineers are investigating and engaging with our supplier. Further updates will be provided when available.
  1. Aiden Spencer says:

    Engineers have completed troubleshooting and have identified a hardware failure. New hardware is en-route to site and will be installed upon arrival, expected within the next 2 hours.

  2. Aiden Spencer says:

    The issue has been resolved and the services have now been restored.

Incident: Virgin Media Circuits

Posted: March 5th, 2020 at 07:50 by Jonathan Clarke

We are currently seeing a number of Virgin Media circuits down. Suppliers have advised that this due to cable damage-causing wider network issues.

Suppliers have advised that engineers are at the location and fibre splicing is taking place.

Further updates will be provided on an hourly basis.

  1. Jonathan Clarke says:

    Suppliers have advised that two cables have been damaged, 96f and 160f.

    Current Situation:
    Work has commenced on the 160f to restore critical services.
    A cable gang is due onsite by 8:30 am with 2km of cable to replace the existing 96f.

    Next steps:
    New 96f cable to be pulled and spliced.

    Further updates to follow.

  2. david labouchardiere says:

    The supplier advises that work continues on the 160f to restore critical services(blue light etc)
    The original 96f still remains unavailable. We have received confirmation that the cable gang has arrived on-site and have begun their prep work to replace the existing 96f.

    Next steps
    New 96f cable to be pulled and spliced. A splicer has now arrived on-site to replace the existing splicer with further splicing resource still expected to arrive at 11am .

    Further updates to follow

  3. david labouchardiere says:

    The supplier had updated to advise that work continues on the 160f to restore services. They have seen a large number of these services come back online but there is still further work to be done on the remaining fibre tubes on this 160F Fibre.
    The original 96f still remains unavailable. The cable gang have been working on this issue all morning with splice work now beginning on the impacted 96F.

    We will continue to provide updates once available

  4. david labouchardiere says:

    The supplier advises that resplicing work is still underway with engineers continuing to work to restore services. We will continue to provide updates once they are made available

  5. david labouchardiere says:

    The supplier has advised that the fibre repair work has been completed. Any circuits still offline should try rebooting their equipment before contacting the service desk for further assistance

Incident: DSL authentication

Posted: February 7th, 2020 at 11:23 by david labouchardiere

We are currently investigating a problem with authentication on DSL sessions. The issue will be preventing any new sessions from establishing but any circuits that are currently connected will remain online unless they are disconnected or rebooted.

We will provide further updates as soon as they are available

  1. david labouchardiere says:

    We have had confirmation that this issue is now resolved. Please ask any affected customers to reboot their router and contact the service desk should they have any further problems.

    We apologise for any inconvenience caused by this incident.

Emergency Maintenance: GLO4-rtr-1.core

Posted: January 29th, 2020 at 18:17 by Adam Heath

Thursday 30th January 2020, 03:00 – 04:00

During the above window we will be performing emergency maintenance on the above device under recommendation of our vendor after errors are being seen.  All circuits connected to this device will experience a service disruption whilst a reload of a card takes place.  We are not expecting the disruption to service to last any more than 10-15 minutes, however, the window is required so that any necessary pre and post change checks can be performed.

We will update upon completion.

  1. Aiden Spencer says:

    This has now been completed

Incident: Kirklees

Posted: January 28th, 2020 at 11:00 by Daniel Jones

We are currently aware of an issue affecting circuits within the Kirklees area. There are ongoing investigations in relation to this issue and we will keep you informed of progress.

  1. Daniel Jones says:

    Services have now been restored. We are still investigating the root cause.

Incident: VoIP calls

Posted: January 23rd, 2020 at 11:43 by david labouchardiere

We have been made aware that one of our VoIP upstream carriers has seen a drop in call volumes and has raised an incident to investigate further. This will affect calls on hosted and sip only accounts. We will provide further updates once they become available

  1. david labouchardiere says:

    Our upstream carrier has implemented a fix in relation to this incident and are monitoring to confirm if the issue is resolved.

Planned Maintenance: pwanstack.wolv6.edge

Posted: January 15th, 2020 at 14:16 by Jamie Smith
Thursday 30th January 2020 00:00 – 03:00 

During the above window we will be carrying out service affecting maintenance on our PWAN edge device located in our Wolverhampton Data Center.

We expect all services terminated on this device to have a maximum downtime of 3 hours. Services include PWAN firewalls, PWAN co-location and psuedowires terminated into customer PWAN solutions.

  1. Jamie Smith says:

    This work is now starting

  2. Jamie Smith says:

    Work has now been completed.

Incident: Edinburgh

Posted: January 14th, 2020 at 09:45 by David Brewis
We are currently aware of an issue affecting circuits within the Edinburgh area. There are ongoing investigations in relation to this issue and we will keep you informed of progress.
  1. David Brewis says:

    Investigations highlight this only impacted leased line services in Edinburgh. We are starting to see services restore as we continue investigations to find the root cause.

  2. David Brewis says:

    After a state of monitoring all services restored and remained stable by 10:15. We apologise for any inconvenience this caused.